OpenStudy (anonymous):

Please help Stefan flips a coin three times. Which of the following shows the number of possible outcomes? 2 6 8 9

jimthompson5910 (jim_thompson5910):

There are 2^3 = 2*2*2 = 8 possible outcomes

OpenStudy (anonymous):

thanks